OpenStudy (anonymous):

find the angle of intersection at (4,-3,2) of spheres x^2+y^2+z^2=29 and x^2+y^2+z^2+4x-6y-8z-47=0.

OpenStudy (anonymous):

Let's find gradient x^2+y^2+z^2+4x-6y-8z-47=0. {4 + 2 x, -6 + 2 y, -8 + 2 z} {12,-12,-4} {2 x, 2 y, 2 z} {8,-6,4} let's plug in number \[\frac{(\{12,-12,-4\}.\{8,-6,4\})}{8 \sqrt{551}}\text{==}\text{Cos}[a]\] \[\text{ArcCos}\left[\sqrt{\frac{19}{29}}\right]\]=0.62762

OpenStudy (anonymous):

approximately 35 degree

OpenStudy (anonymous):

thanks